# Reflection of Nichols Algebras over Coquasi-Hopf Algebras

### v1

@@ Abstract
This paper extends the foundational reflection theory of Nichols algebras to the setting of some certain coquasi-Hopf algebras. Our primary motivation arises from the classification of pointed finite-dimensional coquasi-Hopf algebras. We develop a reflection theory for tuples of simple Yetter-Drinfeld modules in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a 3-cocycle on $G$. We prove that such a tuple gives rise to a semi-Cartan graph if admitting all reflections. Consequently, its Weyl groupoid is well-defined. We further establish several criteria for the finite-dimensionality of Nichols algebras in terms of the associated semi-Cartan graph. As an application, we provide a new proof for the infinite-dimensionality of a specific class of Nichols algebras previously studied in huang2024classification, bypassing extensive computational arguments.
@@ 1 Introduction
The classification of finite-dimensional pointed Hopf algebras has been one of the central problems in the theory of Hopf algebras over the past three decades. A crucial breakthrough in this direction was achieved through the systematic study of Nichols algebras. Heckenberger introduced the revolutionary concepts of semi-Cartan graphs and Weyl groupoids [Hec06, Hec09, Hec10], which provide powerful combinatorial tools for understanding the structure and classification of Nichols algebras.
The reflection theory of Nichols algebras represents one of the most important developments in the field. The fundamental theorem states that a tuple of simple Yetter-Drinfeld modules over Hopf algebra with bijective antipode will give rise to a semi-Cartan graph provided that it admits all reflections [AHS10,HS10]. This result was proved again using a categorical approach in the setting of Hopf algebras in HSdualpair.
@@
We denote by $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ to be the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
@@
In the context of Nichols algebras, the Weyl groupoid captures the symmetries and reflection properties of the underlying braided vector space, establishing a bridge between the algebraic structure of Nichols algebras and the combinatorics of root systems [AHS10].
@@
The interplay between Weyl groupoid theory and reflection theory has led to spectacular progress in the classification of finite-dimensional Nichols algebras. The classification program, initiated by Andruskiewitsch and Schneider in a series of papers [AS98,AS00,AS02,AS10], aims to determine all finite-dimensional pointed Hopf algebras over algebraically closed fields of characteristic zero whose group of group-like elements is abelian. A cornerstone of this classification is Heckenberger’s complete classification of finite-dimensional Nichols algebras of diagonal type [Hec09]. Building on this classification, Angiono made a breakthrough contribution by providing explicit presentations by generators and relations for all Nichols algebras of diagonal type [A13, A15]. The lifting problem of Nichols algebra of diagonal type has been addressed through cocycle deformation [AA+14, AG19, AA17]. The classification of finite-dimensional Nichols algebras over non-abelian groups has achieved excellent progress as well [AGn99,AGn03,AHS10,MS00,FK99,Gn00,Gn+11,HLV12,HS10,HS10b,HS10c,HS14,Hs15,HV17].
@@
In this paper, we extend the reflection theory to the setting of some special coquasi-Hopf algebras. Our motivation comes from the classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras, which has been studied in [rank1,coquasitriangular,tame-rep,FQQR2,QQG,huang2024classification,LiLiu1]. In the classification of pointed Hopf algebras over abelian groups, Nichols algebras of non-diagonal type do not appear. However, in the coquasi-Hopf algebra setting, Nichols algebras of non-diagonal type do occur and present greater challenges. The authors in huang2024classification used extensive computations to establish that a certain class of such Nichols algebras is infinite-dimensional. We later provided an alternative proof of their main result via categorical Morita equivalence LiLiu1. Nonetheless, these arguments are not entirely natural, and we hope that the theory of semi-Cartan graphs can be extended to coquasi-Hopf algebras. If such a generalization holds, it could be applied to the classification of finite-dimensional coquasi-Hopf algebras over non-abelian groups.
@@
Our approach combines the methods of Simon and HSdualpair together. The next theorem generalizes [HSdualpair, Theorem 7.1] to the case of coquasi-Hopf algebras with bijective antipodes under some minor conditions, which is crucial to description of reflection of Nichols algebras.
@@
Now we restrict to the case of pointed cosemisimple coquasi-Hopf algebras. Let $G$ be a finite group, and $\Phi$ be a $3$-cocycle on $G$, which gives rise to the Yetter-Drinfeld module category $^{G}_{G}\mathcal{YD}^{\Phi}$} QQG. Let $\theta\geq 1$ be a positive integer, $\mathbb{I}=\{1,2,..,\theta\}$ and $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ are irreducible Yetter-Drinfeld modules. We always denote $\mathcal{B}(M)$ to be the Nichols algebra of $M_{1}\oplus\cdots\oplus M_{\theta}$ in the following context. Suppose $M$ admits $i$-th reflection for some $1\leq i\leq\theta$ (see Definition 5.1). Then the $i$-th reflection of $M$ is denoted by $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, which is a tuple consisting of finite-dimensional irreducible Yetter-Drinfeld modules. There is a natural explanation of reflection of Nichols algebra in terms of $\Omega$:
@@
(Theorem 5.7) With above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$, and $\mathcal{B}(M_{i})$ is finite-dimensional. Then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
Readers who are familiar with reflection theory in the Hopf algebra context will recognize this formula, we emphasize that our contribution lies in establishing this result in the framework of coquasi-Hopf algebras. The extension from Hopf algebras to coquasi-Hopf algebras introduces technical challenges, particularly in sections 3 and 4 where the non-strict monoidal structure necessitates a reworking of the reflection theory. Now with the help of above theorems, our main result follows naturally.
@@
After establishing that $M$ gives rise to a semi-Cartan graph, several additional criteria for determining the finite-dimensionality of the Nichols algebra can also be derived. We show that if $\mathcal{B}(M)$ is finite-dimensional, then the semi-Cartan graph $\mathcal{G}(M)$ is a finite semi-Cartan graph. Furthermore, if $\mathcal{G}(M)$ is moreover a standard finite semi-Cartan graph, that is $A^{X}=A^{Y}$, for all $X,Y\in\mathcal{X}$, then $A^{M}$ must be a finite Cartan matrix. As an application of these results, we provide a third proof of [huang2024classification, Proposition 4.1].
@@ 2.1 Coquasi-Hopf algebras
By definition, coquasi-Hopf algebras are exactly the dual of Drinfeld’s quasi-Hopf algebras Drinfeld. Their formal definition can be given as follows.
@@
  • A convolution-invertible map $\Phi:H^{\otimes 3}\rightarrow\mathbbm{k}$ called associator,
@@
Following the definition in [preantipode], we now recall the definition of a preantipode, which plays an important role in further calculations.
@@
[preantipode, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
@@
Furthermore, [preantipode, Remark 3.7] provides us a useful property:
@@
Meanwhile, when $H$ is equipped with the structure of a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then preantipode for $H$ always exists.
@@
[preantipode, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then
@@ 2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [YD-module]. It is defined as the center $\mathcal{Z}({{{}^{H}}\mathscr{M}})$ of the comodule category $^{H}\mathscr{M}$}. Hence it is braided monoidal equivalent to $\textbf{Rep}(D(H))$. Since it plays a crucial role in our paper, we recall this definition below.
@@
[YD-module, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
@@
$\overset{(\ref{2.18})}{=}s(\mathcal{S}(h_{1}),h_{3}v_{-1})t(h_{2},v_{-2})v_{-3}\otimes(h_{4}\rhd v_{0})$
$\overset{(\ref{2.17})}{=}s(\mathcal{S}(h_{1}),h_{5}v_{-1})t(h_{3},v_{-3})\mathcal{S}(h_{2})(h_{4}v_{-2})\otimes h_{6}\rhd v_{0}$
@@
For our purpose, we aim to describe Yetter-Drinfeld modules over coquasi-Hopf algebra of the form $(\mathbbm{k}G,\Phi)$, where $G$ is a finite group and $\Phi$ a $3$-cocycle on $G$. Assume that $V$ is a left $kG$-comodule with comodule structure map $\delta_{L}:V\rightarrow kG\otimes V$. Define
@@
In general, let $\mathcal{C}$ be a braided monoidal category and $A$ a Hopf algebra in $\mathcal{C}$. The category of Yetter-Drinfeld modules $^{A}_{A}\mathcal{YD}(\mathcal{C})$} is well-defined. The reader may refer to [rootsys, Section 3] for details.
@@ 2.3 Bosonization for coquasi-Hopf algebras
Bosonization for Hopf algebras is well-known. In [quasibon], the authors introduced bosonization for quasi-Hopf algebras. Dually, bosonization for coquasi-Hopf algebras is given in YD-module. Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. Suppose $R$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}, we denote
@@
[YD-module, Theorem 5.2] Let us consider on $N:=R\otimes H$ the following structures:
@@
Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
@@
[YD-module, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we denote
@@
For reader’s convenience, we still give an example about $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$. Suppose $R$ is a Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Since $R$ is a left $G$-comodule, there is a $G$-grading on $R$:
@@
where $^{x}R:=\{X\in R\mid\rho(x)=x\otimes X\}$. For simplicity, in this paper, homogeneous elements in $^{G}_{G}\mathcal{YD}^{\Phi}$} are denoted by capital letters, say $X,Y,Z$..., and the associated degrees are denoted by their lower cases, say $x,y,z$
@@
The definition of the bosonization naturally extends to the case that $R$ is a Hopf algebra in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from B95 and AF00 that will be essential for the construction presented in Section 3.
@@
[AF00, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$. Let $\pi:H\to A$ and $\iota:A\to H$ be Hopf algebra morphisms such that $\pi\circ\iota=\mathrm{id}_{A}$. If $\mathcal{C}$ has equalizers and $A\otimes(-)$ preserves equalizers, there is a Hopf algebra $R$ in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, such that
@@
[B95, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}. There is an obvious isomorphism of braided monoidal categories
@@
2.4 Nichols algebras in braided monoidal category

### v2

# arXiv:2512.04560v2[math.QA] 27 Feb 2026

@@ Abstract
This paper extends the foundational reflection theory of Nichols algebras to the setting of certain coquasi-Hopf algebras. Our primary motivation arises from the classification of pointed finite-dimensional coquasi-Hopf algebras. We develop a reflection theory for tuples of simple Yetter-Drinfeld modules in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a 3-cocycle on $G$. We prove that such a tuple gives rise to a semi-Cartan graph if it admits all reflections. Consequently, its Weyl groupoid is well-defined. We further establish several criteria for the finite-dimensionality of Nichols algebras in terms of the associated semi-Cartan graph. As an application, we provide a new proof for the infinite-dimensionality of a specific class of Nichols algebras previously studied in Huang et al. , bypassing extensive computational arguments.
Mathematics Subject Classification: 16T05; 18M15; 20F55
@@ 1 Introduction
The classification of finite-dimensional pointed Hopf algebras has been one of the central problems in the theory of Hopf algebras over the past three decades. A crucial breakthrough in this direction was achieved through the systematic study of Nichols algebras. Heckenberger introduced the revolutionary concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for understanding the structure and classification of Nichols algebras.
The reflection theory of Nichols algebras represents one of the most important developments in the field. The fundamental theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode will give rise to a semi-Cartan graph provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b]. This result was proved again using a categorical approach in the setting of Hopf algebras in Heckenberger and Schneider (2013a).
@@
Let $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ denote the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
@@
In the context of Nichols algebras, the Weyl groupoid captures the symmetries and reflection properties of the underlying braided vector space, establishing a bridge between the algebraic structure of Nichols algebras and the combinatorics of root systems [Andruskiewitsch et al., 2010].
@@
The interplay between Weyl groupoid theory and reflection theory has led to spectacular progress in the classification of finite-dimensional Nichols algebras. The classification program, initiated by Andruskiewitsch and Schneider in a series of papers [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2002,Andruskiewitsch and Schneider, 2010], aims to determine all finite-dimensional pointed Hopf algebras over algebraically closed fields of characteristic zero whose group of group-like elements is abelian. A cornerstone of this classification is Heckenberger’s complete classification of finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2009]. Building on this classification, Angiono made a breakthrough contribution by providing explicit presentations by generators and relations for all Nichols algebras of diagonal type [Angiono, 2013,Angiono, 2015]. The lifting problem of Nichols algebras of diagonal type has been addressed through cocycle deformation [Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017]. The classification of finite-dimensional Nichols algebras over non-abelian groups has achieved excellent progress as well [Andruskiewitsch and Graña, 1999,Andruskiewitsch and Graña, 2003,Andruskiewitsch et al., 2010,Milinski and Schneider, 2000,Fomin and Kirillov, 1999,Graña, 2000,Graña et al., 2011,Dello and Zhang, 2016,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013b,Heckenberger and Schneider, 2010a,Heckenberger and Vendramin, 2014,Heckenberger and Vendramin, 2015,Heckenberger and Vendramin, 2017].
@@
In this paper, we extend the reflection theory to the setting of some special coquasi-Hopf algebras. Our motivation comes from the classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras, which has been studied in [Huang et al., 2011,Huang and Liu, 2012,Huang et al., 2015,Huang et al., 2021,Huang et al., 2020,Huang et al., ,Li and Liu, 2025]. In the classification of pointed Hopf algebras over abelian groups, Nichols algebras of non-diagonal type do not appear. However, in the coquasi-Hopf algebra setting, Nichols algebras of non-diagonal type do occur and present greater challenges. The authors in Huang et al. used extensive computations to establish that a certain class of such Nichols algebras is infinite-dimensional. We later provided an alternative proof of their main result via categorical Morita equivalence Li and Liu (2025). Nonetheless, these arguments are not entirely natural, and we hope that the theory of semi-Cartan graphs can be extended to coquasi-Hopf algebras. If such a generalization holds, it could be applied to the classification of finite-dimensional coquasi-Hopf algebras over non-abelian groups.
@@
Our approach combines the methods of Barvels et al. (2015) and Heckenberger and Schneider (2013a). The next theorem generalizes [Heckenberger and Schneider, 2013a, Theorem 7.1] to the case of coquasi-Hopf algebras with bijective antipodes under some minor conditions, which is crucial to the description of the reflection of Nichols algebras.
@@
Now we restrict to the case of pointed cosemisimple coquasi-Hopf algebras. Let $G$ be a finite group, and $\Phi$ be a $3$-cocycle on $G$, which gives rise to the Yetter-Drinfeld module category $^{G}_{G}\mathcal{YD}^{\Phi}$} Huang et al. (2020). Let $\theta\geq 1$ be a positive integer, $\mathbb{I}=\{1,2,..,\theta\}$ and $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{G}_{G}\mathcal{YD}^{\Phi}}$ are irreducible Yetter-Drinfeld modules. We always denote $\mathcal{B}(M)$ as the Nichols algebra of $M_{1}\oplus\cdots\oplus M_{\theta}$ in the following context. Suppose $M$ admits the $i$-th reflection for some $1\leq i\leq\theta$ (see Definition 5.1). Then the $i$-th reflection of $M$ is denoted by $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, which is a tuple consisting of finite-dimensional irreducible Yetter-Drinfeld modules. There is a natural explanation of reflection of Nichols algebra in terms of $\Omega$:
@@
(Theorem 5.7) Under the above assumptions on M, suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$, and $\mathcal{B}(M_{i})$ is finite-dimensional. Then there is an isomorphism of Hopf algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}:
@@
Readers who are familiar with reflection theory in the Hopf algebra context will recognize this formula, we emphasize that our contribution lies in establishing this result in the framework of coquasi-Hopf algebras. The extension from Hopf algebras to coquasi-Hopf algebras introduces technical challenges, particularly in Sections 3 and 4 where the non-strict monoidal structure necessitates a reworking of the reflection theory.with the help of the above theorems, our main result follows naturally.
@@
After establishing that $M$ gives rise to a semi-Cartan graph, several additional criteria for determining the finite-dimensionality of the Nichols algebra can also be derived. We show that if $\mathcal{B}(M)$ is finite-dimensional, then the semi-Cartan graph $\mathcal{G}(M)$ is a finite semi-Cartan graph. Furthermore, if $\mathcal{G}(M)$ is a standard finite semi-Cartan graph, that is $A^{X}=A^{Y}$, for all $X,Y\in\mathcal{X}$, then $A^{M}$ must be a finite Cartan matrix. As an application of these results, we provide a third proof of [Huang et al., , Proposition 4.1].
@@ 2.1 Coquasi-Hopf algebras
By definition, coquasi-Hopf algebras are exactly the dual of Drinfeld’s quasi-Hopf algebras Drinfeld (1989). Their formal definition can be given as follows.
@@
  • A convolution-invertible map $\Phi:H^{\otimes 3}\rightarrow\mathbbm{k}$ called an associator,
@@
Following the definition in [Ardizzoni and Pavarin, 2012], we now recall the definition of a preantipode, plays an important role in subsequent calculations.
@@
[Ardizzoni and Pavarin, 2012, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
@@
Furthermore, [Ardizzoni and Pavarin, 2012, Remark 3.7] provides us a useful property:
@@
Meanwhile, when $H$ is equipped with the structure of a coquasi-Hopf algebra with an antipode $(\mathcal{S},\alpha,\beta)$, then preantipode for $H$ always exists.
@@
[Ardizzoni and Pavarin, 2012, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with an antipode $(\mathcal{S},\alpha,\beta)$, then
@@ 2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of the Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [Ardizzoni and Pavarin, 2013]. It is defined as the center $\mathcal{Z}({{{}^{H}}\mathscr{M}})$ of the comodule category $^{H}\mathscr{M}$}. Hence it is braided monoidal equivalent to $\textbf{Rep}(D(H))$. Since it plays a crucial role in our paper, we recall this definition below.
@@
[Ardizzoni and Pavarin, 2013, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
@@
[…]
@@
For our purpose, we aim to describe Yetter-Drinfeld modules over coquasi-Hopf algebra of the form $(\mathbbm{k}G,\Phi)$, where $G$ is a finite group and $\Phi$ a $3$-cocycle on $G$. Assume that $V$ is a left $kG$-comodule with a comodule structure map $\delta_{L}:V\rightarrow kG\otimes V$. Define
@@
In general, let $\mathcal{C}$ be a braided monoidal category and $A$ a Hopf algebra in $\mathcal{C}$. The category of Yetter-Drinfeld modules $^{A}_{A}\mathcal{YD}(\mathcal{C})$} is well-defined. The reader may refer to [Heckenberger and Schneider, [2020] ©2020, Section 3] for details.
@@ 2.3 Bosonization for coquasi-Hopf algebras
Bosonization for Hopf algebras is well-known. In [Bulacu and Nauwelaerts, 2002], the authors introduced bosonization for quasi-Hopf algebras. Dually, bosonization for coquasi-Hopf algebras is given in Ardizzoni and Pavarin (2013). Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. Suppose $R$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}, we denote
@@
[Ardizzoni and Pavarin, 2013, Theorem 5.2] Let us consider on $N:=R\otimes H$ the following structures:
@@
Now suppose $N$ and $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
@@
[Ardizzoni and Pavarin, 2013, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we denote
@@
For the reader’s convenience, we provide an example about $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite group and $\Phi$ is a $3$-cocycle on $G$. Suppose $R$ is a Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Since $R$ is a left $G$-comodule, there is a $G$-grading on $R$:
@@
where $^{x}R:=\{X\in R\mid\rho(x)=x\otimes X\}$. For simplicity, in this paper, homogeneous elements in $^{G}_{G}\mathcal{YD}^{\Phi}$} are denoted by capital letters, say $X,Y,Z$..., and the associated degrees are denoted by their lowercase letters, say $x,y,z$
@@
The definition of the bosonization naturally extends to the case that $R$ is a Hopf algebra in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from Bespalov (1995) and Alonso Alvarez and Fernández Vilaboa (2000) that will be essential for the construction presented in Section 3.
@@
[Alonso Alvarez and Fernández Vilaboa, 2000, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$. Let $\pi:H\to A$ and $\iota:A\to H$ be Hopf algebra morphisms such that $\pi\circ\iota=\mathrm{id}_{A}$. If $\mathcal{C}$ has equalizers and $A\otimes(-)$ preserves equalizers, there is a Hopf algebra $R$ in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, such that
@@
[Bespalov, 1995, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}. There is an obvious isomorphism of braided monoidal categories
@@
2.4 Nichols algebras in a braided monoidal category